Groups St Andrews 1997 in Bath

Geoff Smith, Treasurer of Groups St Andrews 1997, School of Mathematical Sciences, University of Bath, Bath BA2 7AY, England

Conference email address: groups97@dcs.st-and.ac.uk

The conference arrival day is Saturday 26th July 1997. The conference departure day is Saturday 9th August 1997. Sunday 3rd August is a rest day. We gratefully acknowledge the financial support of the Edinburgh and London Mathematical Societies.

The general format will be similar to previous Groups St Andrews meetings. We have five speakers giving courses of lectures in the first week, and a wide diversity of talks and invited lectures in the second week. The lecture courses will be given by Laszlo Babai (Chicago), Martin Bridson (Oxford), Chris Brookes (Cambridge), Cheryl Praeger (Western Australia) and Aner Shalev (Jerusalem).

In the second week there will be two theme days celebrating mathematics in the spirit of William Burnside and Roger Lyndon. The programmes on these days will be co-ordinated by Efim Zelmanov (Yale) and Charles F. Miller III (Melbourne) respectively.

Local conference arrangements are all being handled by Geoff Smith at Bath (G.C.Smith@bath.ac.uk). Anything to do with money or accommodation is his province.

Talks and Papers

A form is enclosed which enables you to request time to present a talk at the meeting. Colin Campbell and Edmund Robertson are dealing with this aspect of the organization, so this form should be returned to St Andrews. The proceedings of the conference will be published by Cambridge University Press as usual, and articles will only appear in the proceedings subject to a reviewing process.

Pre-Registration and Registration Fees

Please find attached a Pre-Registration Form which you can use to reserve accommodation. The arrangements are designed to be very economical, to encourage post-graduate participation and to make the conference accessible to as many people as possible. In particular, substantial discounts are available in return for early payment. We encourage everyone to pre-register and register as early as possible to avoid disappointment. We have accommodation reserved to cope with numbers who attended Galway in 1993. If there is more demand, later bookings may have to be turned away.

We appreciate that people based in countries with no banking infrastructure may find it difficult to arrange for payment before the conference and in time to qualify for discounts. At the discretion of the Treasurer, the discounts may remain available to such persons up until July 26th 1997 (though they should return the various forms, and firmly commit themselves to coming).

A small non-returnable pre-registration fee is payable by all persons intending to attend Groups St Andrews 1997 in Bath, and by all accompanying persons. This fee is UK £25 or US $45, but is reduced to UK £15 or US $27 provided payment is received in Bath on or before March 15th 1997.

The conference Registration Fee is UK £140 (US $252) payable on or before arrival on July 26th 1997. However, there is a substantial discount available and the registration fee drops to UK £110 (US$198) if payment is received in Bath on or before May 31st 1997. This fee includes the cost of the conference proceedings, coffee and tea, and the social events.

There will be an attractive social programme, and non-mathematical accompanying adults are welcome to participate fully. This includes the conference dinner, a civic reception, coach trips and the conference cricket match.

Accompanying persons aged 18 years or over pay full pre-registration fees and half registration fees (with proportionate discounts for prompt payment). They are not entitled to tea and coffee except by special arrangement with the Treasurer.

Children aged 17 or younger pay the pre-registration fee but are exempt from the registration fee. Children may go on the conference afternoon coach trips, but may not attend other conference events unless special arrangements are made with the Treasurer.

All payments may be made only in one of the following forms.

1. Cash in sterling (British currency).

2. Cash in United States dollars.

3. A sterling cheque drawn on a British bank.

4. A bankers draft in sterling (British currency).

The fourth form of payment is the obvious method for people based outside the United Kingdom.

Cheques should be made payable to Groups 1997. Paper receipts will be available at the conference, and email receipts will be sent automatically provided you supply an email address.

Accommodation

There are four accommodation options. All campus accommodation is self-catering, and people on a tight budget therefore have the option of preparing all their own meals. Bookings may be made for the first week (leaving Saturday August 2nd) or the second week (arriving Sunday 3rd August) or both. Booking is by the week, and if you book for both weeks, you (of course) also get the use of your room on the night of Saturday August 2nd.

Option A Single room, £100 per week (discounted to £80 per week if payment is received in Bath by May 31st 1997).

Option B Single room, £130 per week (discounted to £112 per week if payment is received in Bath by May 31st 1997).

Option C Single room, £160 per week (discounted to £140 per week if payment is received in Bath by May 31st 1997).

Option D Private flat for four, £270 per week (discounted to £240 per week if payment is received in Bath by May 31st 1997).

Payment is also acceptable in US dollars (cash only) at the rate of UK £1 = US$1:80.

Descriptions: All accommodation is at least of a reasonable standard.

Option A Washbasin in room. Shared toilets and showers.

Option B Ensuite facilities (shower, toilet and washbasin).

Option C Full spacious ensuite facilities (including bath).

Option D A self contained flat for four people. We have only a limited number of such flats available. This is a particularly attractive option for families with small children (extra camp beds can be supplied). It would be sensible to return your Pre-Registration Form immediately if you wish to try to reserve a flat.

Catering

The University of Bath’s main catering facilities will be occupied by an Open University conference throughout Groups St Andrews 1997 in Bath. However, there are a couple of food outlets on campus which will sell rolls and sandwiches, and remain open during the evening. We suggest that the following regime may prove healthy and attractive.

Breakfast: You have this in your room or kitchen, purchasing food and drink for this purpose from the campus shop or in Bath city centre.

Lunch: We expect that most people will not want to have to prepare their own lunches during the limited break in the conference programme in the middle of the day. We have therefore arranged for good quality ‘meals on trays’ to be delivered to the accommodation area at lunchtime every day except for the rest day on Sunday August 3rd. The menus will vary widely, but will avoid foods likely to give rise to health concerns. A vegetarian option will be available (but it must be ordered in advance). Persons with specific dietary requirements (medical or religious) should write to the Treasurer as soon as possible and he will try to help.

These meals will be delivered only to people who have ordered them, and it is therefore imperative that you apply for these lunches in time, should you wish to take them.

The price is £42 for 7 lunches during Sunday July 27th - Saturday August 2nd.

Sunday August 3rd is a rest day and we assume that people will make their own arrangements for lunch.

The price is £30 for 5 lunches during Monday August 4th - Friday August 8th.

Dinner: Conferees are responsible for finding their own dinner. One evening there will be a conference dinner off campus which may be attended by everyone who has paid a registration fee (even half fee). On other evenings you may wish to eat in your own kitchen, or to take a short bus trip to the centre of Bath. The city contains a very large number of restaurants specializing in diverse cuisines. Many Public Houses serve inexpensive evening meals. A comprehensive list of restaurants, cafés, take-aways and public houses will be available on arrival, together with expert reviews.

The buses run every few minutes between campus and the city centre. The journey time is about 10 minutes. Bath is a small, relaxed, friendly town, and most people find it is a very safe place.

Refreshments

The Senior Common Room bar of the University of Bath will be available to conferees every evening during the conference. At this time of year very few of the academic staff of the University use this facility in the evenings, so we expect to have (almost) exclusive use of the bar.

Smoking

The University of Bath operates a no smoking policy save for certain designated areas. There is such a designated area in a room close to the SCR bar. Almost all British restaurants and public houses permit smoking, and a list of the few exceptions in Bath will be available on arrival.

Summary

Early booking and early payment is advised to ensure the accommodation of your choice.

1. Send the Pre-Registration Form, with payment, to Bath now – or at least to arrive by March 15th 1997 otherwise the price goes up. State on the form what sort of accommodation you require.

2. Send the Registration Form with your registration fee and accommodation fee to arrive by May 31st 1997 ( otherwise the price goes up and there may not be space for you!).

3. If you want to give a talk, fill in the Talk Form and send it to arrive in St Andrews before May 31st 1997 (otherwise there is little chance that you will get a slot).

4. Send the Lunch Form with payment to arrive before July 1st 1997 (if you require lunch).

5. Payments can only be made using one of the four specified methods. It is acceptable to lump together payments – for example by writing a UK cheque in sterling to cover lunches, registration and accommodation.

6. Web users should peruse the conference web site http://www.bath.ac.uk/∼masgcs/gps97/ where more and more information about the conference programme, the city of Bath, travel advice, social events and so on will appear as time goes by. This is in addition to information which will be sent out in electronic or paper form to all who pre-register.

Pre-Registration Form

Groups St Andrews 1997 in Bath For discounted rates, form and payment must arrive by 15th March 1997

Name: Address for written correspondence:

Email and fax addresses if possible: Telephone numbers if possible: Preferred badge name and affiliation (including title if desired): Names of accompanying persons: My first preference for accommodation is A/B/C/D. My second preference for accommodation is A/B/C/D. I seek accommodation for week 1 / week 2 / both weeks / neither week (please specify). I provisionally expect that I and my party will/will not want lunch delivered (delete as necessary). A non-returnable pre-registration fee is payable by all persons intending to attend Groups St Andrews 1997 in Bath, and by all accompanying persons. This fee is UK £25 or US $45, but is reduced to UK £15 or US $27 provided payment is received in Bath on or before March 15th 1997.

• I enclose payment for the appropriate number of pre-registrationfees in the form of:

(Recall the restrictions on methods of payment; cheques should be made in favour of Groups 1997.) Space for any personal comments or requests (e.g. I request a formal written invitation to attend the conference):

Please return this form to: Dr Geoff Smith Groups St Andrews 1997 in Bath School of Mathematical Sciences University of Bath, Bath BA2 7AY, England fax: +44 1225 826492

Lunch Form

Groups St Andrews 1997 in Bath

This form must arrive in Bath before July 1st 1997

Name:

Affiliation:

Recall that the prices are UK £42 for 7 lunches during Sunday July 27th - Saturday August 2nd and UK £30 for 5 lunches during Monday August 4th - Friday August 8th.

I request lunches in week 1 / week 2 / both weeks / not at all (delete as appropriate).

Explain if you are paying for more than one person, and if you are vegetarian or have other dietary requirements.

• I enclose appropriate payment in the form of (insert details):

(Recall the restrictions on methods of payment; cheques should be made in favour of Groups 1997. The dollar exchange rate is UK £1 = US$1:80 as usual.)

Please return this form to: Dr Geoff Smith Groups St Andrews 1997 in Bath School of Mathematical Sciences University of Bath Bath BA2 7AY, England fax: +44 1225 826492

Registration Form

Groups St Andrews 1997 in Bath

For discounted rates, form and payment must arrive by May 31st 1997

Name:

Address for written correspondence:

Names of accompanying persons:

My first preference for accommodation is A/B/C/D.

My second preference for accommodation is A/B/C/D.

• I enclose appropriate payment for my first preference accommodation in the form of (insert details):

(Recall the restrictions on methods of payment; cheques should be made in favour of Groups 1997. If you cannot be assigned accommodation of the type you specify as a first preference, a financial correction must and will be made at the conference.)

• I enclose appropriate registration fee(s) in the form of (insert details):

The conference Registration Fee is UK £140 (US $252) payable on or before arrival on July 26th 1997. However, there is a substantial discount available and the registration fee drops to UK £110 (US $198) if payment is received in Bath on or before May 31st 1997.

Please return this form to: Dr Geoff Smith Groups St Andrews 1997 in Bath School of Mathematical Sciences University of Bath Bath BA2 7AY, England fax: +44 1225 826492

Talk Form

This form must arrive in St Andrews before May 31st 1997

Name:

Affiliation:

Address for written correspondence:

Email and Fax address (if possible):

I wish to give a talk at Groups St Andrews 1997 in Bath entitled:

I request a minimum of (insert number of minutes):

I wish to submit a paper associated with this talk for consideration for inclusion in the Conference Proceedings Yes/No (delete as appropriate). Survey articles are particularly welcome. Where possible the typesetting should be in plain L^{A}TEX. The use of symbol definitions and macros should be avoided in order to simplify editorial work. Electronic versions of the articles may be sent to groups97@dcs.st-and.ac.uk

Burnside day will be on Monday August 4th and Lyndon day will be on Thursday August 7th. Please indicate if you would like to apply to speak on one of these theme days, and if so, which one.

Please return this form to:

The Scientific Committee Groups St Andrews 1997 in Bath Mathematical Institute University of St Andrews North Haugh St Andrews Fife KY16 9SS Scotland fax +44 1334-463748